OpenStudy (anonymous):

Solve for x in terms of a. log3(x + a) − log3(x − a) = 1

OpenStudy (anonymous):

compress: log((x+a/(x-a))=1 then write in exponential form 10=(x+a)/(x-a)

OpenStudy (anonymous):

cross multiply: 10x-10a=x+a 9x=11a x=11a/9

OpenStudy (anonymous):

is it \[\log_{3}(x+a) \]?

OpenStudy (anonymous):

according to tku's problem the logarithm has a base of 3

OpenStudy (anonymous):

i thought it was a coefficient, not a base. i wasn't sure because of how he wrote it.

OpenStudy (anonymous):

it s base

OpenStudy (anonymous):

ok so then: 3=(x+a)/(x-a) 3x-3a=x+a 2x=4a x=2a